OpenStudy (anonymous):

Ethyl acetate reacts with H2, in the presence of a catalyst to yield ethyl alcohol: C4H8O2 + 2H2--> 2C2H6O. A. How many grams of ethyl alcohol are produced by reaction of 2.7 mol of ethyl acetate with H2? B. How many grams of ethyl alcohol are produced by reaction of 13.0g of ethyl acetate with H2? C. How many grams of H2 are needed to react with 13.0g of ethyl acetate?

OpenStudy (abb0t):

Well, start with what you're given first. You have 2.7 mol EtOAc, and you want EtOH, which is 2 mol. So, start by setting up the conversion (this is called, \(dimensional ~analysis\)) So, you have: \(\sf \color{blue}{2.7~mol~EtOAc \times \frac{2~mol~EtOH}{1~mol~EtOAc}}\) Now, you should notice that the \(\sf \color{red}{mol~EtOAc}\) cancel out! Do you follow me so far? Your next step is to molar mass! Remember that to find molar mass, which has units of: \(\sf \color{}{\frac{mass}{mol}=\frac{g}{mol}}\), now you're basically doing same thing as above, you cancel out mol to get grams! That answers part \(B\). \(\sf \color{purple}{NOTE:}\) TO FIND Molar mass, \(\sf \color{orange}{\sum n(atom)}\), which is the number of atoms given, for instance, you have up above: C\(_2\), so the mass of that would be [2 \(\times\) (mass of 1 carbon)] + [sum of hydrogen atom]+...

OpenStudy (anonymous):

Thank you very much for your help!